### v2

# Retrieving Implicit and Explicit Emotional Events Using Large Language Models

Abstract
Large language models (LLMs) have garnered significant attention in recent years due to their impressive performance. While considerable research has evaluated these models from various perspectives, the extent to which LLMs can perform implicit and explicit emotion retrieval remains largely unexplored. To address this gap, this study investigates LLMs’ emotion retrieval capabilities in commonsense. Through extensive experiments involving multiple models, we systematically evaluate the ability of LLMs on emotion retrieval. Specifically, we propose a supervised contrastive probing method to verify LLMs’ performance for implicit and explicit emotion retrieval, as well as the diversity of the emotional events they retrieve. The results offer valuable insights into the strengths and limitations of LLMs in handling emotion retrieval.
[…]
Retrieving Implicit and Explicit Emotional Events Using Large Language Models
Guimin Hu♠, Hasti Seifi† ♠University of Copenhagen, ‡Arizona State University, United States hasti.seifi@asu.edu, rice.hu.x@gmail.com
1 Introduction
Large language models (LLM) have drawn considerable interest in recent years due to their remarkable performance. In the constantly evolving realm of artificial intelligence, Large Language Models (LLMs) have risen as cutting-edge tools for natural language understanding. In community of sentiment analysis, LLMs’ emotion retrieval abilities come under scrutiny of the commonsense retrieval domain. Past research has delved into the potential of language models for emotion recognition Hu et al. (2024), emotion cause analysis Zheng et al. (2022); Hu et al. (2021) and sentiment analysis Hu et al. (2022b), as seen in studies Lee and Lee (2021); Hu et al. (2022a); Xia and Ding (2019).
| Emotion | Emotional Event | Explicit | Implicit |
| Joy | be proud of someone | ✗ | ✓ |
| Joy | successful career | ✗ | ✓ |
| Joy | glad someone helped | ✓ | ✗ |
| Sad | feel lonely | ✗ | ✓ |
| Sad | I feel like a loser | ✗ | ✓ |
| Sad | sad because they didn’t win | ✓ | ✗ |
| Angry | angry at not being able to control oneself | ✓ | ✗ |
| Angry | feel angry at someone’s laziness | ✓ | ✗ |
| Angry | he lost his temper | ✗ | ✓ |
Generally speaking, receiving praise typically elicits a feeling of happiness in the individual. On the contrary, criticism often leads to feelings of pessimism and frustration. We refer to this type of event, which can trigger a specific emotion in humans, as an “emotional event”. From this perspective, emotional reasoning overlaps with common-sense reasoning in certain situations (for example, most people feel happy when praised, which can be considered common sense). Previous research works on analysizing sentiment, detecting emotion category and extracting emotion cause, demonstrating LLM capacity on emotion related tasks. However, these investigations take LLMs as backbone, train a model with supervised learning, and then evaluate the model performance on recognizing emotion for given text or extract emotion cause based on emotion cause, which focuses on correlations from data, ignoring the exploration of LLMs’ own ability on implicit and explicit emotion retrieval within the scope of common sense. Therefore, in this study, we aim to fill this gap by providing both qualitative and quantitative analyses of LLM’s ability to reason implicit and explicit emotion in common sense. Our goal is to evaluate the models’ performance in retrieving about emotional events across various emotion categories, both implicit and explicit, with a focus on addressing the following three questions:
RQ1: Assessment of LLMs on emotion events retrieval across different emotion categories
To address this question, we evaluate the capacity of LLMs on emotion retrieval via taking specific emotion as input to retrieving emotional events those are relevant to the given emotion (e.g., “Joy”, “Sad”, and “Angry”).
RQ2: What are the differences in retrieval performance between implicit and explicit emotional events for various LLMs?
To answer this question, we evaluate how accurately LLMs retrieve implicit and explicit emotional events, where implicit emotional event denote that there are no explicit sentiment words about the topic.
RQ3: How does the diversity of retrieved emotional events vary across LLMs at different levels of precision?
To address this, we employ the supervised contrastive probe to investigate the retrieval ability of LLM and conduct experiments to evaluate the abilities of LLM’s retrieving diverse emotional events.
In summary, our contributions to tackle these challenges are: (1) we conduct a comprehensive evaluation of state-of-the-art Large Language Models (LLMs) in emotional event retrieval, with a specific focus on implicit and explicit emotion retrieval, using public commonsense knowledge base, (2) our analysis is thorough and multifaceted, encompassing both qualitative and quantitative aspects into the strengths and limitations of LLMs in emotion retrieval of commonsense, and (3) we introduce a novel supervised contrastive probe and this probe is tested on its ability to generalize to objects that were not seen during training.
2 Dataset
We employ the emotion-cause flow in Li et al. (2022) as the evaluation benchmark of emotion commonsense retrieval in LLMs. Li et al. (2022) is Chinese commonsense knowledge base which contains concept flow, event flow, emotion intention flow and emotion-cause flow. We collect emotional events corresponding to specific emotion from based on emotion-cause flow. For emotion-cause flow, each emotion category is corresponding to multiple emotional events. Meanwhile, we manually annotate the emotional events as explicit or implicit. Table 1 illustrates some typical cases of and Table 2 shows the details of data split.
| Train | Valid | Test | Total | Explicit | |
| joy | 1922 | 247 | 540 | 2704 | 1305 |
| sad | 370 | 64 | 256 | 1245 | 171 |
| angry | 851 | 138 | 102 | 541 | 173 |
| total | 3143 | 449 | 898 | 4481 | 1649 |
3 Method
3.1 Problem Formalization
Given an emotion category, our objective is to assess the ability of a large language model (LLM) to retrieve emotional events of specified emotion. We frame this as a retrieval task and evaluate the LLM’s performance on emotion event retrieval from the following perspectives: a) the precision of retrieved emotional events on specific emotion, b) the precision of retrieved emotional events on implicit and explicit emotion and c) the diversity of retrieved emotional events.
3.2 A Supervised Contrastive Probe
Given an emotion category and its corresponding emotional event , we construct the pair consisting of emotion and its corresponding emotional event as the objective of contrastive learning, where emotion is formalized into “the event that describe ”. Note that each emotional event corresponds to one emotion category (e.g., “successful career brings someone joy”). We view the emotion category as the label of the pair of emotion and emotional event. For a target example, it has multiple positive samples and multiple negative samples, where positive sample refers to the emotional events that are labels by the emotion same as the target sample, negative sample denotes the pairs those are with the labels differs from target’s label. In the training phrase, we use the standard contrastive loss objective:
where Suppose we partition into mutually exclusive train and test sets, . If we learn via minimizing training loss and these transformations generalize to , i.e., for :
We thus use precision@K over , where a prediction as counted as correct of the correct label is in the set of top K most similar objects, to evaluate the precision of retrieved emotional events given specific emotion.
4 Experiments
4.1 Setup
Hyperparameters
We conducted a light grid search over the hyperparameters, specifically testing the learning rate and the temperature coefficient . We used a batch size of 64 and trained for 15 epochs. Crucially, the optimal hyperparameters (and the best epoch for early stopping) were determined using a validation set randomly sampled from the training data. That is, none of the hyperparameters were tuned based on held-out performance on .
Precision and Diversity
Given an emotion category, LLMs retrieved a set of events corresponding to the emotion. We adopt Precision@K and Diversity@K (D@K) to evaluate the model performance, where Precision@K (P@K) counted as correct of the correct label is in the set of highest K scores and Diversity@K (D@K) denotes the fraction of retrieved unique emotional events in the set of highest K that are relevant to the given emotion. The metrics are given as follows:
where denotes the number of relevant emotional events in the set of highest K scores, denotes the number of all retrieval results, and denotes the number of de-duplicated emotional events in the set of highest K scores.
4.2 Main Results
Figure 1 shows the performance comparison of different Large Language Models (LLMs) in terms of precision at three levels: P@3, P@10, and P@50, for retrieving emotional events across three categories: Joy, Sadness, and Anger. “Joy” consistently has the highest precision across all models, with MPT and MISTRAL being the top performers. Precision drops sharply for emotions like “Sad” and “Angry” across all models, indicating that emotion retrieval for these emotions is more challenging. MPT consistently shows the highest performance for “Joy” in all settings, while MISTRAL also performs well, though slightly behind MPT.
The results suggest that some models like MPT and MISTRAL perform better for detecting “Joy” but struggle with other emotions, particularly “Angry”. Models like GPT show consistently poor performance across the board, indicating weaker “Angry” events retrieval capabilities.
| Model | Diversity@K | Joy(%) | Sad(%) | Angry(%) |
| BERT | D@3 | 0.56 (3) | 0.8 (2) | 3.5 (4) |
| D@10 | 1.9 (10) | 0.8 (2) | 9.8 (11) | |
| D@50 | 13.5 (72) | 17.39 (44) | 80.35 (90) | |
| T5 | D@3 | 0.56 (3) | 0.8 (2) | 0.89 (1) |
| D@10 | 1.9 (10) | 0.8 (2) | 0.0 (0) | |
| D@50 | 7.31 (39) | 14.22 (36) | 16.96 (19) | |
| GPT | D@3 | 0.56 (3) | 1.58 (4) | 2.6 (3) |
| D@10 | 1.9 (10) | 0.4 (1) | 9.8 (11) | |
| D@50 | 9.3 (50) | 0.4 (1) | 67.85 (76) | |
| MPT | D@3 | 0.75 (4) | 1.5 (4) | 2.6 (3) |
| D@10 | 2.1 (11) | 4.3 (11) | 9.8 (10) | |
| D@50 | 10.31 (55) | 50.19 (127) | 44.64 (50) | |
| Mega | D@3 | 0.94 (9) | 1.58 (4) | 3.5 (4) |
| D@10 | 2.1 (11) | 0.4 (1) | 9.8 (11) | |
| D@50 | 10.31 (55) | 0.4 (1) | 73.21 (87) | |
| Llama | D@3 | 0.75 (4) | 1.58 (4) | 3.5 (4) |
| D@10 | 2.1 (11) | 4.3 (11) | 9.8 (11) | |
| D@50 | 10.69 (57) | 66.79 (169) | 88.39 (99) | |
| Mistral | D@3 | 0.75 (4) | 1.58 (4) | 3.5 (4) |
| D@10 | 2.1 (11) | 4.3 (11) | 9.8 (11) | |
| D@50 | 11.06 (59) | 67.19 (170) | 89.28 (92) |
4.3 Implicit VS Explicit Emotion Retrieval
The table 4 presents the retrieval rate of explicit emotions for various models (BERT, T5, GPT, etc.) across three emotional categories: Joy, Sadness, and Anger at different precision levels (P@3, P@10, P@50). T5 and Mistral perform well for Joy at high precision, especially at P@3 and P@10. Anger retrieval is consistent across models, with moderate performance. Sad retrieval is the weakest for all models, with most failing to retrieve explicit Sadness at smaller precision levels, though some improvement is seen at P@50. Additionally, Joy retrieval tends to be the strongest across most models, especially at P@3 and P@10. Models like T5 and Mistral consistently perform well in identifying joyful emotions, although their performance drops as the retrieval set size increases (P@50). Sad retrieval is consistently poor across the models, with some improvements at P@50 for models like T5 and BERT, but generally, most models struggle to retrieve explicit Sad emotions.
| Model | P@K | Joy(%) | Sad(%) | Angry(%) |
| BERT | P@3 | 50.23 | 0.0 | 0.0 |
| P@10 | 76.93 | 0.0 | 39.78 | |
| P@50 | 53.10 | 20.33 | 29.23 | |
| T5 | P@3 | 96.14 | 0.0 | 0.0 |
| P@10 | 75.11 | 0.0 | 0.0 | |
| P@50 | 38.93 | 47.15 | 5.78 | |
| GPT | P@3 | 66.67 | 0.0 | 35.05 |
| P@10 | 70 | 0.0 | 35.05 | |
| P@50 | 70 | 0.0 | 35.05 | |
| MPT | P@3 | 66.67 | 0.0 | 35.05 |
| P@10 | 70 | 0.0 | 35.05 | |
| P@50 | 48.51 | 13.69 | 30 | |
| Mega | P@3 | 66.67 | 0.0 | 36.42 |
| P@10 | 70 | 0.0 | 36.42 | |
| P@50 | 50 | 0.0 | 29.50 | |
| Llama | P@3 | 30.53 | 0.0 | 0.0 |
| P@10 | 76.18 | 8.73 | 33.5 | |
| P@50 | 48.80 | 0.0 | 28.38 | |
| Mistral | P@3 | 54.35 | 0.0 | 0.0 |
| P@10 | 72.78 | 9.82 | 34.37 | |
| P@50 | 48.40 | 13.62 | 27.61 |
4.4 Diversity of Retrieved Emotion Events after De-duplication
We compare the capacity of LLM on retrieving diverse events for given emotion, as shown in Table 4. 3. For each model, we report the results on P@3, P@10 and P@50 for Joy, Sad and Angry. We can observe that most models perform not well on diverse emotional event retrieval. Mistral and Mega show the best overall performance at high diversity levels, particularly excelling in retrieving unique angry events. BERT performs well for Anger at D@50 but lags behind in Joy and Sadness. Models like T5 and GPT demonstrate weaker performance in capturing the diversity of emotions, especially for Sadness.
5 Related Work
Procrustes Analysis
is a statistical technique used to compare shapes or configurations of data by transforming them in a way that minimizes the differences between corresponding points. It is often applied in fields like shape analysis, psychometrics, and computer vision. The core idea is to align two datasets as closely as possible through transformations such as translation, scaling, and rotation, without changing the fundamental structure of the shapes. Procrustes Analysis as a probing tool has been applied in concept alignment Peng and Søgaard (2024) and cross-modal alignment Ma et al. (2022); Li et al. (2023). For example, Peng et al. Peng and Søgaard (2024) evaluated concept alignment on multilingual LLMs by revisiting the traditional bilingual dictionary induction task. Li et al. Li et al. (2023) presented an empirical evaluation across four families of LMs (BERT, GPT-2, OPT, and LLaMA-2) and three vision model architectures (ResNet, SegFormer, and MAE) on concept space alignment.
6 Conclusion
In this study, we evaluated the performance of various Large Language Models (LLMs) in retrieving emotional events across three categories: Joy, Sadness, and Anger, at different precision levels (P@3, P@10, and P@50). The results indicate that models exhibit distinct strengths and weaknesses depending on the emotion and dataset size. The performance differences across models and emotions highlight the need for further improvement in emotion recognition tasks, particularly for underrepresented emotions like Sadness. Future work should focus on developing more balanced models that can handle a broader range of emotional content with higher precision across all categories.
Limitation
The experiments may rely on a specific dataset or a narrow set of emotional scenarios, which could limit the generalizability of the findings. The performance of these models might vary significantly when tested on more diverse datasets or with different forms of emotional expressions (e.g., multimodal inputs like text, audio, or video).